OpenStudy (anonymous):

Find the function rule. X Y -2 9 -1 4 0 -1 1 -6 2 -11 Please help I am sooooooo confused.

OpenStudy (anonymous):

Try plotting these points on a graph and look at the shape of the curve, if it's a straight line, your function would be \[f(x)=mx+c\] if it's parabolic it could be something like \[f(x)=\frac{x^2}{4a}\] or\[f(x)=\sqrt{4ax}\] See what kind of graph you're getting